a)    Music Class

For the host IP address, identify the prefix

192.168.20.10/??

Identify the LSB (least significant bit)

Identify the Subnet ID to which the host belongs

 

Identify the broadcast address

 

Identify the usable range of host IP addresses

 


1
Expert's answer
2021-07-07T01:52:26-0400

For 50 hosts, block size which fulfills the host requirement "=64= 2^6"

For block size 64, number of hosts bits "=2^6 = 6" hosts bits (least significant bits)

For 5 hosts bits, number of network bits "=32-6 =26" network bits

When network bit is 26, sub-net mask =26 =8|8|8|2 = 255.255.255.92


When network IP is 192.168.20.0 and block size is 64 then broadcast IP = (192.168.20.0+64) - 1 = 192.168.20.63


When network IP is 192.168.20.0 and broadcast IP is 192.168.20.63 then IP range = 192.168.20.1 - 192.168.20.62 (usable IP starts just after network IP and ends just before broadcast IP)